# Predictive Coding Networks and Inference Learning: Tutorial and Survey

### v1

@@ 2.3.2. Learning Algorithms
Table 3 shows the key steps involved in BP and IL. Conceptually, the main difference is the forward ($1\rightarrow L$) and backward ($L\rightarrow 1$) sweeps of BP, absent in IL. In contrast, updates in IL (for both activations and weights) use only local information. Updates are local in time: all layers could in principle be updated in parallel (see fig. 6). In contrast, in BP, weight updates in lower layers cannot be done before computations in all higher layers have been completed. BP thus involves waiting times; it is not local in time. If this parallelization of computation can be realized, this difference implies a potential speed-up of IL compared to BP (see appendix C). It should be noted that the learning time in practice involves several additional factors, discussed further in section 4.
A final note on learning concerns weight initialization. It is well-known in machine learning that in deep networks, correctly initializing weights is required for networks to be trainable (goodfellow_deep_2016, ; prince_understanding_2023, ; schoenholz2017deep; Bukva:2023ksv). This has not yet been studied in relation to PC except (frieder_non-convergence_2022, ), which only showed that weights ought to be sufficiently small at initialization. Practical uses so far (e.g., (song_inferring_2024, )) and publicly available implementations (tschantz_infer-activelypypc_2023, ) use a value of 0.05, similar to typical values chosen for ANNs (paszke_pytorch_2019, ), but this choice lacks a theoretical basis, and more optimal initialization schemes may exist.
Figure 6. Locality in IL. Schematic llustration of weight updates BP in an FNN vs. IL in a PCN. In the former, the cost or loss function $\mathcal{L}$ plays the role of the (squared) error in this example, which must be propagated from the output layer all the way back through the network until the relevant weight is reached; in contrast, in IL the errors are locally computed. This is often cited as a reason for the greater biological plausibility of PCNs, but is practically relevant insofar as it allows the updates to be performed in parallel, and avoids issues surrounding vanishing/exploding gradients. Figure adapted from (millidge_predictive_2022, ).
@@ 2.3.3. Structure
A typical, perhaps universal representation of the structure of an FNN is a diagram like in fig. 4. The connections in this diagram are unambiguously defined by the activity rule (1). The learning algorithm (BP) is defined by (3) and (4) (typically not included in visual representations). In this way, FNNs admit some degree of independence between the notions of structure and learning algorithm.
The story becomes more complicated in PCNs. Unlike FNNs, they do not (yet) have a universally employed visual representation as FNNs fig. 4. Most diagrams shows that PCNs have bottom-up connections from data to labels (predictions), as well as top-down connections from labels to data (errors), meaning they have feedback/recurrence [^5] It is also worth emphasizing that this recurrence is of an altogether different nature than that in recurrent neural networks (RNNs) (goodfellow_deep_2016, ), where the recurrence is in time rather than in spatial connectivity (i.e., RNNs are characterized by weight sharing). In PCNs, time steps are iterations in the inference phase, whereas in RNNs, time steps are subsequent samples of temporal data. A variation of PCNs applied to temporal data is (millidge_predictive_2024, ). in all activation nodes. This is a fundamental feature of PC. Note however, that (for discriminative PCNs) these feedback components are only used during training (cf. section 2.2.5).
What, then, precisely defines the structure of a PCN? PCNs use activity nodes like FNNs, and adds error nodes through (8); but to get connections between layers like in FNNs, one should also include predictions, (7). However, illustrations like fig. 5 also include feedback connections, a key aspect of PCNs – hence they could additionally be included under structure. Mathematically, one way of defining these connections is by the activity rule (10), meaning this equation could also be included under structure. However, this equation is derived from the minimum of the objective function, which also defines the learning algorithm. This illustrates how structure and learning algorithm are entangled in PCNs.
This discussion explains why the term predictive coding in the literature sometimes appears to relate to a learning algorithm, and sometimes to different network architectures. [^6] Note that this ambiguity of the term predictive coding is further complicated by applications of the term to altogether different models, e.g. deep learning models with an architecture that includes separate top-down and bottom-up models, which are both trained with BP. Indeed, we have explained that it is in fact both. Nonetheless, some aspects of these algorithms are clearly ‘structural’ in nature: hence we shall keep using this terms when useful, as e.g. in fig. 2.
@@ 2.3.4. Computational Complexity
Since testing a FNN and PCN is equivalent, their time complexity during this phase is equal. As for the training phase, by studying the computations in table 3, the time complexity of a single weight update can be computed (reviewed in detail in appendix C). The results of these computations are summarized in table 4. With matrix multiplications being the most costly computations and defining $M$ as the complexity for the largest weight multiplication, the complexity can be shown to be $\mathcal{O}(LM)$ for BP, $\mathcal{O}(TLM)$ for IL (salvatori_stable_2023, ; alonso_understanding_2023) and $\mathcal{O}(LM)$ for Incremental IL (see also appendix C). That is, BP and incremental IL have the same time complexity per weight update, with standard IL a factor of $T$ more costly.
Table 4. Computational Complexity for BP, IL, and incremental IL/incremental PC. $M$ is the most costly matrix computation, $T$ is the number of steps to convergence, and $L$ is the number of layers (i.e., depth). As discussed above, the local nature of IL allows parallelization to remove the dependence on depth, which is not possible in BP.
Standard Parallelized
BP $\mathcal{O}(LM)$ $\mathcal{O}(LM)$
IL $\mathcal{O}(TLM)$ $\mathcal{O}(TM)$
Incremental IL $\mathcal{O}(LM)$ $\mathcal{O}(M)$
However, as mentioned above, the computations in IL enjoy both temporal and spatial locality, meaning the computations in different layers during inference and learning could be parallelized. Ignoring overhead, this implies that IL’s time complexity would decrease to $\mathcal{O}(TM)$, and incremental IL would have $\mathcal{O}(M)$, which is faster than BP. In this case, total training time no longer scales with the depth of the network, which is a highly desirable feature. A first implementation of this kind was provided by (salvatori_stable_2023, ). However, their algorithm included a substantial computational overhead, such that it is only faster in networks where $L\gtrsim 128$. At the same time, BP is heavily optimized by dedicated libraries that employ GPU acceleration, whereas no similarly comprehensive library yet exists for IL. In addition, note that the time complexity per weight update is not the only factor that determines training time in practice. The second key factor is the number of epochs until convergence, which depends on the optimizer, dataset, and other hyperparameters used. These can impact the learning properties of BP and IL in different ways. We discuss some of these topics in section 4.
@@ 2.3.5. Empirical results
As is well known to deep learning practitioners, many factors affect performance of FNNs, which is no different with discriminative PCNs. Exponential growth of the search space with hyperparameters makes a systematic comparison intractable, meaning existing works compare FNNs and PCNs only over a small range of hyperparameters. Interestingly, discriminative PCNs often perform extremely similar to FNNs (whittington_approximation_2017, ; song_inferring_2024, ), but works by (frieder_non-convergence_2022, ; frieder_bad_2024, ) also make clear that care needs to be taken in importing knowledge and intuitions from BP-trained networks since IL does introduce meaningful changes related to its mathematical properties. With this remark we summarize the advantages and limitations that have been observed so far.
In terms of improvements of PCNs over FNNs, most noteworthy are those observed in (song_inferring_2024, ) and (alonso_theoretical_2022, ). Ref. (song_inferring_2024, ) observes minor but statistically significant improvements for online learning (learning with batch size 1) and data efficiency (learning with less than 300 data points per class), on the order of 2% in the best case. Gains in classification accuracy of this order is also observed in (alonso_theoretical_2022, ), who additionally observe that the use of optimizer greatly affects results, exhibiting differences between BP and IL. Specifically, IL combined with SGD sometimes converges to poor local minima, while BP does not (with Adam, IL again improves to the level of BP). Somewhat more impressive are the improvements observed for continual learning tasks and concept drift tasks observed in (song_inferring_2024, ), with gains of 5-20%. Overall, both works generally observe faster convergence measured as number of epochs. The improvement is typically small but sometimes substantial. The recent work (innocenti_understanding_2023, ) finds evidence confirming this; in section 4, we discuss theoretical insights that help explain these results. On larger architectures (e.g., for CNNs), (alonso_theoretical_2022, ) observe that IL with SGD sometimes converges to shallow local minima, such that BP outperforms IL by a large margin (order of 10%). However, this difference is decreased to 2% by using Adam.
Although IL often appears to converge in fewer epochs, its total computation time (as measured by wall-clock time) is often cited to be larger than BP, although few works have studied this extensively. Reasons for this include standard IL’s increased computation per weight update (improved with incremental IL), and an increased computational overhead compared to heavily optimized BP libraries. For standard IL, then, three main counteracting ‘forces’ appear to be at play: (1) decreased number of epochs required, but (2) increased computation per weight update, and (3) increased overhead. With (2) largely removed in incremental IL, it remains to be studied how (1) and (3) balance out in practice.
@@ 2.4. Extensions
It is straightforward to extend discriminative PCNs to more complex architectures such as convolutional neural networks (salvatori_predictive_2021, ; alonso_theoretical_2022, ), recurrent neural networks (salvatori_predictive_2021, ), variational autoencoders and transformers (pinchetti_predictive_2022, ). Other objective functions may also be considered (pinchetti_predictive_2022, ; millidge_theoretical_2022, ).
@@ 2.4.1. Convolutional Layers
Convolutional layers make the replacement
(15) $a^{\ell+1}_{i}=f\Big{(}\underbrace{\sum_{j}W_{ij}^{\ell}a^{\ell}_{j}+b^{\ell}_{j}}_{z^{\ell}_{i}}\Big{)}\rightarrow a^{\ell+1}_{ij}=f\Big{(}\underbrace{\sum_{x=1}^{k_{\ell}}\sum_{y=1}^{k_{\ell}}W_{xy}^{\ell}a_{i+x,\,j+y}^{\ell}+b^{\ell}}_{z^{\ell}_{ij}}\Big{)}$
where $W^{\ell}$ is the kernel, a matrix of size $k_{\ell}\times k_{\ell}$ with $k_{\ell}$ the kernel size. For each (now two-dimensional) output neuron $a^{\ell+1}_{ij}$ one sums over $x$ and $y$, the neurons in the local receptive field. $W$, then, is a weight matrix shared by all neurons, together with a single bias $b$. Layer-dependent stride $s_{\ell}$ (which allows moving the local receptive field in larger steps) and padding of zeros $p_{\ell}$ can be introduced by setting $x\rightarrow s_{\ell}x-p_{\ell}$ and $y\rightarrow s_{\ell}y-p_{\ell}$. Importing this to PCNs, we see that we need only change the pre-activations $\bm{z}^{\ell}$ in the relations $\bm{a}^{\ell}=f(\bm{z}^{\ell})$ (ANNs) and $\bm{\mu}^{\ell}=f(\bm{z}^{\ell})$ (PCNs) respectively. Pooling layers (which reduce the dimensionality of layers by downsampling) can similarly be written by changing only $\bm{z}^{\ell}$.
@@ 2.4.2. Other Objective Functions, Transformers and VAEs
So far, the objective function considered was $E=\sum_{\ell}(\bm{\epsilon}^{\ell})^{2}$, which corresponded to the MSE loss in FNNs, ${\mathcal{L}_{n}=\left(\hat{\bm{y}}-\bm{y}^{(n)}\right)^{2}}$. Generalizing this to any layer-dependent energy one may write $E=\sum_{\ell}E_{\ell}$, where $E_{\ell}$ may now be generalized to other functions. For instance, instead of an MSE-like error,(pinchetti_predictive_2022, ) derives cross-entropy-like error:
@@
which is appropriate for binary output variables. This can be understood and generalized further by understanding PCNs as a probabilistic latent variable model, cf. section 3. Even more generally, one may write:
${E}_{\ell}=g(a^{\ell},a^{\ell-1})$
where $g$ is some function (millidge_theoretical_2022, ).
In (pinchetti_predictive_2022, ), PC is extended to train transformers and VAEs. Training transformers requires attention layers which include the softmax function. These require summing over all nodes in a layer, and are hence not included in PC as presented above, which considered only layer-dependent activations. Considering PC for non- Gaussian distributions (cf. section 3) however (pinchetti_predictive_2022, ) makes this extension possible. Using a similar framework, the same authors also train variational autoencoders (VAEs) (kingma_auto-encoding_2022, ) with IL.
@@ 2.5. Generative PCNs
In the supervised learning context, discriminative models approximate the posterior probability distribution of labels given datapoints, $p(y|x)$ (bishop_pattern_2006, ). If equipped with a softmax function at the final layer, this is precisely what the model above does (hence the term discriminative PCN). Making only a minor change to the approach discussed above, one instead obtains a model that approximates $p(x|y)$: a generative model, from which synthetic points in the input space can be generated. With IL, such a model can be used both for supervised learning, as well as unsupervised learning, both of which we discuss below.
In general, it should be noted that testing a generative model is less straightforward than testing discriminative models, since they can be used for several purposes (e.g. density estimation, sampling/generation, latent representation learning) (prince_understanding_2023, ).
@@ 2.5.1. Supervised Learning
Structure. To obtain a generative PCN, one changes the direction of the local predictions in the network: $\bm{\mu}^{\ell}=f(\bm{w}^{\ell-1}\bm{a}^{\ell-1})$ becomes $\bm{\mu}^{\ell}=f(\bm{w}^{\ell+1}\bm{a}^{\ell+1})$. [^7] Equivalently, one can leave the local predictions unchanged, and swap sides where the data and label are clamped: i.e. clamp $\bm{x}^{(n)}$ to $\bm{a}^{L}$ and $\bm{y}^{(n)}$ to $\bm{a}^{0}$, cf. section 2.2.1.
Training Procedure. The training procedure and learning algorithm (IL), are unchanged: (10) is run until converged, followed by a weight update. The changed local prediction results in slight changes to update rules: in (10) and (12), $\ell+1$ becomes $\ell-1$.
Testing Procedure. Clamping the final layer $\bm{a}^{L}$ to a label, with the reversed prediction direction one can derive the reverse of (14): a synthetic datapoint is created at $\bm{a}^{0}$ in a single feedforward pass (without requiring several iterations of (10)).

### v2

@@ 2.3.2. Learning Algorithms
Table 3 shows the key steps involved in BP and IL. The main difference is the forward ($1\rightarrow L$) and backward ($L\rightarrow 1$) sweeps of BP, absent in IL. In contrast, updates in IL (for both activations and weights) use only local information. Updates are local in time: all layers could in principle be updated in parallel (see fig. 6). In contrast, in BP, weight updates in lower layers cannot be done before computations in all higher layers have been completed. BP thus involves waiting times; it is not local in time. If this parallelization of computation can be realized, this difference implies a potential speed-up of IL compared to BP. It should be noted that the learning time in practice involves several additional factors, discussed further in section 4.
A final note on learning concerns weight initialization. It is well-known in machine learning that in deep networks, correctly initializing weights is required for networks to be trainable (Goodfellow et al., 2016; Prince, 2023; Schoenholz et al., 2017; Bukva et al., 2023). This has not yet been studied in relation to PC except (Frieder and Lukasiewicz, 2022), which only showed that weights ought to be sufficiently small at initialization. Practical uses so far (e.g., (Song et al., 2024)) and publicly available implementations (Tschantz and Millidge, 2023) use a value of 0.05, similar to typical values chosen for ANNs (Paszke et al., 2019), but this choice lacks a theoretical basis, and more optimal initialization schemes may exist.
Figure 6. Locality in IL. Schematic llustration of weight updates BP in an FNN vs. IL in a PCN. In the former, the cost or loss function $\mathcal{L}$ plays the role of the (squared) error in this example, which must be propagated from the output layer all the way back through the network until the relevant weight is reached; in contrast, in IL the errors are locally computed. This is often cited as a reason for the greater biological plausibility of PCNs, but is practically relevant insofar as it allows the updates to be performed in parallel, and avoids issues surrounding vanishing/exploding gradients. Figure adapted from (Millidge et al., 2022a).
@@ 2.3.3. Structure
A typical representation of the structure of an FNN is a diagram like in fig. 4. The connections in this diagram are unambiguously defined by the activity rule (1). The learning algorithm (BP) is defined by (3) and (4) (typically not included in visual representations). In this way, FNNs admit some degree of independence between the notions of structure and learning algorithm. In contrast, PCNs do not (yet) have a universally employed visual representation, although most diagrams are similar to fig. 5. Observe that PCNs have bottom-up connections from data to labels (predictions), as well as top-down connections from labels to data (errors), meaning they have feedback/recurrence [^7] It is also worth emphasizing that this recurrence is of an altogether different nature than that in recurrent neural networks (RNNs) (Goodfellow et al., 2016), where the recurrence is in time rather than in spatial connectivity (i.e., RNNs are characterized by weight sharing). In PCNs, time steps are iterations in the inference phase, whereas in RNNs, time steps are subsequent samples of temporal data. A variation of PCNs applied to temporal data is (Millidge et al., 2024). in all nodes, a fundamental feature of PC. Note however, that (for discriminative PCNs) these feedback components are only used during training (cf. section 2.2.4). For PCNs, structure and learning algorithm are more difficult to separate (cf. appendix B.3).
@@ 2.3.4. Computational Complexity
Since testing a FNN and PCN is equivalent, their time complexity during this phase is equal. As for the training phase, by studying the computations in table 3, the time complexity of a single weight update can be computed (reviewed in detail in appendix C). The results of these computations are summarized in table 4. With matrix multiplications being the most costly computations and defining $M$ as the complexity for the largest weight multiplication, the complexity can be shown to be $\mathcal{O}(LM)$ for BP, $\mathcal{O}(TLM)$ for IL and $\mathcal{O}(LM)$ for Incremental IL (Salvatori et al., 2024b; Alonso et al., 2024). That is, BP and incremental IL have the same time complexity per weight update, with standard IL a factor of $T$ more costly.
Table 4. Time Complexity for BP, IL, and incremental IL/incremental PC. $M$ is the most costly matrix computation, $T$ is the number of steps to convergence, and $L$ is the number of layers (i.e., depth). As discussed above, the local nature of IL allows parallelization to remove the dependence on depth, which is not possible in BP.
BP IL Incremental IL
Standard $\mathcal{O}(LM)$ $\mathcal{O}(TLM)$ $\mathcal{O}(LM)$
Parallelized $\mathcal{O}(LM)$ $\mathcal{O}(TM)$ $\mathcal{O}(M)$
However, as mentioned above, the computations in IL enjoy both temporal and spatial locality, meaning the computations in different layers during inference and learning could be parallelized. Ignoring overhead, this implies that IL’s time complexity would decrease to $\mathcal{O}(TM)$, and incremental IL would have $\mathcal{O}(M)$, which is faster than BP. In this case, total training time no longer scales with the depth of the network, which is a highly desirable feature. A first implementation of this kind was provided by (Salvatori et al., 2024b). However, their algorithm included a substantial computational overhead, such that it is only faster in networks where $L\gtrsim 128$. At the same time, BP is heavily optimized by dedicated libraries that employ GPU acceleration, whereas no similarly comprehensive library yet exists for IL. In addition, note that the time complexity per weight update is not the only factor that determines training time in practice: the second key factor is the number of epochs until convergence, which depends on the optimizer, dataset, and other hyperparameters used. These can impact the learning properties of BP and IL in different ways. We discuss some of these topics in section 4.
@@ 2.3.5. Empirical Results
As is well known to deep learning practitioners, many factors affect performance of FNNs, which is no different with discriminative PCNs. Exponential growth of the search space with hyperparameters makes a systematic comparison intractable, meaning existing works compare FNNs and PCNs only over a small range of hyperparameters. Interestingly, discriminative PCNs often perform extremely similar to FNNs (Whittington and Bogacz, 2017; Song et al., 2024), but works by (Frieder and Lukasiewicz, 2022; Frieder et al., 2024) also make clear that care needs to be taken in importing knowledge and intuitions from BP-trained networks since IL does introduce meaningful changes related to its mathematical properties. With this remark we summarize the advantages and limitations that have been observed so far.
In terms of improvements of PCNs over FNNs, most noteworthy are those observed in (Song et al., 2024) and (Alonso et al., 2022). Ref. (Song et al., 2024) observes minor but statistically significant improvements for online learning (learning with batch size 1) and data efficiency (learning with less than 300 data points per class), on the order of 2% in the best case. Gains in classification accuracy of this order is also observed in (Alonso et al., 2022), who additionally observe that the use of optimizer greatly affects results, exhibiting differences between BP and IL. Specifically, IL combined with SGD sometimes converges to poor local minima, while BP does not (with Adam, IL again improves to the level of BP). Somewhat more impressive are the improvements observed for continual learning tasks and concept drift tasks observed in (Song et al., 2024), with gains of 5-20%. Overall, both works generally observe faster convergence measured as number of epochs. The recent work (Innocenti et al., 2023) finds evidence confirming this (cf. section 4, for further theoretical discussion). On larger architectures (e.g., for CNNs), (Alonso et al., 2022) observe that IL with SGD sometimes converges to shallow local minima, such that BP outperforms IL by a large margin (order of 10%). However, this difference is decreased to 2% by using Adam.
Although IL often appears to converge in fewer epochs, its total computation time (as measured by wall-clock time) is often cited to be larger than BP, although few works have studied this extensively. Reasons for this include standard IL’s $\mathcal{O}(TLM)$ complexity per weight update (improved with incremental IL), and an increased computational overhead. For standard IL, then, three main counteracting ‘forces’ appear to be at play: (1) decreased number of epochs required, (2) increased computation per weight update, and (3) increased overhead. With (2) largely removed in incremental IL, it remains to be studied how (1) and (3) balance out in practice.
@@ 2.4. Extensions
It is straightforward to extend discriminative PCNs to more complex architectures such as convolutional neural networks (Salvatori et al., 2021b; Alonso et al., 2022), recurrent neural networks (Salvatori et al., 2021b), variational autoencoders and transformers (Pinchetti et al., 2022). Other objective functions may also be considered (Pinchetti et al., 2022; Millidge et al., 2023b).
@@ 2.4.1. Convolutional Layers
In traditional ANNs, convolutional layers replace (1) by
(15) $a^{\ell+1}_{ij}=f\Big{(}\sum_{x=1}^{k_{\ell}}\sum_{y=1}^{k_{\ell}}W_{xy}^{\ell}a_{i+x,\,j+y}^{\ell}+b^{\ell}\Big{)}$
where $W^{\ell}$ is the kernel, a matrix of size $k_{\ell}\times k_{\ell}$ with $k_{\ell}$ the kernel size. For each (now two-dimensional) output neuron $a^{\ell+1}_{ij}$ one sums over $x$ and $y$, the neurons in the local receptive field. $W$, then, is a weight matrix shared by all neurons, together with a single bias $b^{\ell}$. Layer-dependent stride $s_{\ell}$ (which allows moving the local receptive field in larger steps) and padding of zeros $p_{\ell}$ can be introduced by setting $x\rightarrow s_{\ell}x-p_{\ell}$ and $y\rightarrow s_{\ell}y-p_{\ell}$. Importing this to PCNs, we see that we need only change the arguments of the activation functions: $\bm{z}^{\ell}$ in the relations $\bm{a}^{\ell}=f(\bm{z}^{\ell})$ (ANNs) and $\bm{\mu}^{\ell}=f(\bm{z}^{\ell})$ (PCNs) respectively. Pooling layers (which reduce the dimensionality of layers by downsampling) can similarly be written by changing only $\bm{z}^{\ell}$.
@@ 2.4.2. Other Objective Functions, Transformers and VAEs
So far, the objective function considered was $E=\sum_{\ell}(\bm{\epsilon}^{\ell})^{2}$, which corresponded to the MSE loss in FNNs, ${\mathcal{L}_{n}=\left(\hat{\bm{y}}-\bm{y}^{(n)}\right)^{2}}$. Generalizing this to any layer-dependent energy one may write $E=\sum_{\ell}E_{\ell}$, where $E_{\ell}$ may now be generalized to other functions. For instance, instead of an MSE-like error, (Pinchetti et al., 2022) derives cross-entropy-like error:
@@
which is appropriate for binary output variables. Even more generally, one may write:${E}_{\ell}=g(a^{\ell},a^{\ell-1})$ where $g$ is some function (Millidge et al., 2023b).
In (Pinchetti et al., 2022), this is used to train transformers and VAEs. Training transformers requires attention layers which include the softmax function. These require summing over all nodes in a layer, and are hence not included in PC as presented above, which considered only layer-dependent activations. Considering PC for non- Gaussian distributions (cf. section 3) however (Pinchetti et al., 2022) makes this extension possible. Using a similar framework, the same authors also train variational autoencoders (VAEs) (Kingma and Welling, 2014) with IL.
@@ 2.5. Generative PCNs
In the supervised learning context, discriminative models approximate the posterior probability distribution of labels given datapoints, $p(y|x)$ (Bishop, 2006). If equipped with a softmax function at the final layer, this is precisely what the model above does (hence the term discriminative PCN). Making only a minor change to the approach discussed above, one instead obtains a model that approximates $p(x|y)$: a generative model, from which synthetic points in the input space can be generated. Generative modeling is often done using unsupervised learning, concerned with finding patterns and structure in unlabeled data (Bishop, 2006). This means a change in the problem setup: one no longer have a dataset $\{\bm{x}^{(n)},\bm{y}^{(n)}\}_{n=1}^{N}$, but only $\{\bm{x}^{(n)}\}_{n=1}^{N}$.
To obtain a generative PCN, one changes the structure of the network; specifically, the direction of the local predictions: $\bm{\mu}^{\ell}=f(\bm{w}^{\ell-1}\bm{a}^{\ell-1})$ becomes $\bm{\mu}^{\ell}=f(\bm{w}^{\ell+1}\bm{a}^{\ell+1})$. [^8] Equivalently, one leaves the local predictions unchanged, and swap sides where the data and label are clamped: i.e. clamp $\bm{x}^{(n)}$ to $\bm{a}^{L}$ and $\bm{y}^{(n)}$ to $\bm{a}^{0}$, cf. section 2.2. With IL, such a model can be used both for supervised learning, as well as unsupervised learning, with the difference being the training and testing procedures, discussed below.
In general, it should be noted that testing a generative model is less straightforward than testing discriminative models, since they can be used for several purposes (e.g. density estimation, sampling/generation, latent representation learning) (Prince, 2023).
@@ 2.5.1. Supervised Learning
Compared to discriminative PCNs, the training procedure and learning algorithm (IL), are unchanged: (10) is run until converged, followed by a weight update. The changed local prediction results in slight changes to update rules: in (10) and (12), $\ell+1$ becomes $\ell-1$. For the testing procedure, clamping the final layer $\bm{a}^{L}$ to a label, with the reversed prediction direction one has the reverse of (14): a synthetic datapoint is created at $\bm{a}^{0}$ in a single ‘backward’ pass [^9] In a generative PCN, backward is the direction of predictions (cf. section 2.2), i.e. this backward pass involves the same computations as the forward pass in an FNN. (without requiring several iterations of (10)).